Find the percent gain or percent loss when a car bought for PhP 925 000 is sold for



PhP 850 000.

1
Expert's answer
2021-12-13T07:20:43-0500

Selling price - Buying price = gain/loss

850000 - 925000 = - 75000

Percentage loss = loss divided by buying price × 100

Percentage loss = 75000÷925000×100 = 8.108%